Use the information below to answer questions 3.27 to 3.30.

Autonomous consumption = R1000 Autonomous investment = R5000 Autonomous government spending = R20000 Net exports = R4000 Marginal propensity to consume = 0.7 Tax rate = 25%

 

25. The equilibrium level of income is …

[1] R39900 [2] R30000 [3] R63000 [4] R99900

 

 

26. If the tax rate rises to 30% and government spending rises to R25000, ceteris paribus, what is the new equilibrium income?

[1] R63000 [2] R63400 [3] R70000 [4] R68600



1
Expert's answer
2021-07-28T09:52:01-0400

Solution

"25.) Y=C+I+G+X"

"C=1000+0.7(1-tax)Y"


"Y=1000+0.7(Y-0.25)+5000+20000+4000"


"Y=30000+0.7(0.75Y)"


"Y=R63400"


"26.) Y= 35000+0.7(Y-0.30Y)"


"0.51Y=35000"


"Y=R68600"